OpenStudy (anonymous):

There are 15 fewer than three times as many marching band members as there are color guard members. There are 105 members of the marching band. how many members are in the color guard

OpenStudy (anonymous):

\[3c = b - 15\] \[b = 105\] \[3c = 105 - 15\] \[\frac{ 3c }{ 3 } = \frac{ 90}{ 3 }\] c = 30